Factor and solve. A photograph is going to be enlarged to print. The length is 4/3 of the width. If the area of the print is 192 square centimeters, what are the dimensions of the length and width?
Answer by LinnW(1048)

set w = width
set l = length
l = (4/3)w
area = l * w
area = (4/3)w * w
192 = (4/3)w * w
multiply each side by 3/4
144 = w^2
12 = w
since l = (4/3)w ,
l = (4/3)12
l = 16